

********************************************************************
UARS WEEKLY STATUS REPORT - 24 SEP 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to perform nominally, taking solar data on all
available orbital tracking periods. This week we received about 57% of
each day's data amounting to nearly 391 minutes of solar observing time
per day.
***********
SUSIM
***********
SUSIM was turned off on Sept. 23 for the yaw-around. Otherwise and as
available telemetry time permitted, SUSIM successfully gathered daily
mid (1.1 nm) resolution, MgII, occultation, solar pointing offset,
and continuous monitoring scans.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004109 has been reprocessed and archived at the GSFC DAAC. HEPSA Level 2
data files from 2004110 to 2004119 are processed and waiting for archiving.
MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2004109.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 11 September 2004 (Orbit 71150, GMT Day 255)
through 17 September 2004 (Orbit 71255, GMT Day 261).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 256/1343 71173 POWER ON
The beta angle ranged between the angles of 56.0 degrees to 31.8
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 25 September 2004 (DOY 269).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.96 to 1.27 degrees C. Additional
battery performance data for each battery on 10 September and 17
September 2004 are provided in Appendix D.
The clock error ranged from 12.0 to -13.6 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 06
HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The new release of the CMS code continued to undergo testing this
week. This new release was built to fix the file transfer problem
encountered in the previous build. The preliminary testing showed the
file transfer issue to be resolved, but the build still requires all
other process testing. The goal for completion is the end of
September.
The FOT released the new Eclipse software build to operations on 13
September 2004. This new release took care of several DRs that had
been written against the previous release. No major issues have been
found and the software has run well since its installation.
MEETINGS
-None this report period
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 25 September 2004 (DOY 269).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 17 September 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
71155 255/0806 04-258 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
71173 256/1343 04-259 HALOE POWER ON
71200 258/0756 04-260 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
71220 259/1633 04-261 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
71226 260/0111 04-262 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
71242 261/0249 04-263 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
71247 261/1147 04-264 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 17 September 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 17 September 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
71155 255/0809 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
71200 258/0817 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
71220 259/1636 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
71226 260/0131 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
71242 261/0249 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
71247 261/1152 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 17 September 2004
10 September 2004 - GMT Day 254 Beta = 58.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 -11.2/-067.2 +3.87/+2.60 +22.0/-14.4 26.4 1.03 15.9
3 -22.4/-112.0 +2.92/+1.66 +21.6/-14.4 26.2 1.02 15.9
17 September 2004 - GMT Day 261 Beta = 31.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 -00.0/-078.4 +5.14/+3.87 +27.2/-14.4 25.4 1.03 19.5
3 -11.2/-140.0 +3.87/+2.91 +26.4/-14.4 25.4 1.03 19.2
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 17 September 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-060 FOT 71177 256/1904 36935 Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-060 FOT 71177 256/1904 36935 Late Acquisition